### v1

# Corner modes in Non-Hermitian long-range model

Abstract
We consider non-Hermitian (NH) analog of a second-order topological insulator, protected by chiral symmetry, in the presence of second-nearest neighbor hopping elements to theoretically investigate the interplay between long-range and topological order away from Hermiticity. In addition to the four zero-energy corner modes present in the first nearest neighbor hopping model, we uncover that the second nearest neighbor hopping introduces another topological phase with sixteen zero-energy corner modes. Importantly, the NH effects are manifested in altering the Hermitian phase boundaries for both the models. While comparing the complex energy spectrum under open boundary conditions, and bi-orthogonalized quadrupolar winding number (QWN) in real space, we resolve the apparent anomaly in the bulk boundary correspondence of the NH system as compared to the Hermitian counterpart by incorporating the effect of non-Bloch form of momentum into the mass term. The above invariant is also capable of capturing the phase boundaries between the two different topological phases where the degeneracy of the corner modes is evident, as exclusively observed for the second nearest neighbor model.
I Introduction
The systems with topological band properties are identified with gapless boundary modes that are characterized by symmetry-protected topological invariants. This is known as bulk boundary correspondence (BBC) Hasan and Kane (2010); Sato and Ando (2017). The conventional BBC, being an integral part of the first-order ($n=1$) topological phase Hasan and Kane (2010); Haldane (1988); Kane and Mele (2005); Bernevig et al. (2006), is generalized for higher-order ($n>1$) topological phases in $d\geq 2$ dimensions where there exist $n_{c}=(d-n)$-dimensional boundary modes Benalcazar et al. (2017a, b); Song et al. (2017); Langbehn et al. (2017); Schindler et al. (2018); Franca et al. (2018); Wang et al. (2019); Călugăru et al. (2019); Szumniak et al. (2020); Ni et al. (2020); Xie et al. (2021); Saha et al. (2022); Zhu (2018); Liu et al. (2018); Yan et al. (2018); Wang et al. (2018a); Zhang et al. (2019a, b); Volpez et al. (2019); Yan (2019a); Ghorashi et al. (2019); Wu et al. (2020); Laubscher et al. (2020); Roy (2020); Zhang et al. (2020); Kheirkhah et al. (2021); Yan (2019b); Ahn and Yang (2020); Luo et al. (2021); Wang et al. (2018b); Ghosh et al. (2021); Roy and Juričić (2021); Li et al. (2021); Ghosh et al. (2022a, b). For example, the second-order topological insulator (SOTI) hosts zero-dimensional (0D) localized corner modes at zero-energy while this phase is characterized by nested polarization or quadrupolar moment. Very recently, it has been reported that the number of boundary modes in a topological phase can be tuned by considering long-range hopping terms DeGottardi et al. (2013); Niu et al. (2012); Xie et al. (2019); Dias and Marques (2022) as well as by implementing periodic Floquet drive Thakurathi et al. (2013); Ghosh et al. (2022a). Once the extended model continues to preserve the chiral symmetry (CS), one can characterize the new topological phase by winding numbers in odd spatial dimension  Schnyder et al. (2008); Kitaev (2009); Ryu et al. (2010). The number of degenerate zero-energy states at each boundary increases according to the enhancement of the range of the hopping amplitudes as indicated by the winding number ensuring the BBC Ryu and Hatsugai (2002); Teo and Kane (2010). It is noteworthy that one-dimensional winding number for the first-order topological systems becomes passive in case of even-dimensional generalizations Nakahara (2018). In contrast, the higher-order topological (HOT) phase in even spatial dimension can be characterized by an appropriately defined winding number preserving CS as a constraint Benalcazar and Cerjan (2022).
In recent years, thanks to the practical realization of higher-order topological phases in meta-materials Parappurath et al. (2020); Yang et al. (2019); Malzard et al. (2015); Regensburger et al. (2012) where energy conservation no longer holds El-Ganainy et al. (2018); Denner et al. (2021) and the domain of topological quantum matter can be extended to the non-Hermitian (NH) systems. The coupling to the environment  Bergholtz and Budich (2019); Yang et al. (2021); San-Jose et al. (2016), disorder/interaction-mediated quasiparticles with finite lifetime Kozii and Fu ; Yoshida et al. (2018); Shen and Fu (2018) can effectively induce a complex self energy that is modeled by an NH effective Hamiltonian  Musslimani et al. (2008); Makris et al. (2008); Yao et al. (2018); Kawabata et al. (2019); Bergholtz et al. (2021); Sone et al. (2020); Denner et al. (2021); Ghosh and Black-Schaffer . Interestingly, the non-Bloch nature of the wavefunction for the NH systems renormalizes the topological mass term, thus enriching the BBC such that topological phase transitions perceived with open-boundary conditions can be explained by an appropriate bulk invariant  Yao and Wang (2018); Kunst et al. (2018); Liu et al. (2019); Helbig et al. (2020); Borgnia et al. (2020); Koch and Budich (2020); Zirnstein et al. (2021); Takane (2021). The NH topological systems showcase various intriguing features such as the skin effect where the bulk states accumulate at the boundary  Yao and Wang (2018); Kunst et al. (2018); Helbig et al. (2020); Kawabata et al. (2020), exceptional points where eigenstates, corresponding to the degenerate bands, coalesce Bender (2007); Heiss (2012).
Going beyond the scope of short-range hopping, long-range hopping in Hermitian systems is found to mediate versatile topological phases where the number of zero-energy modes increases Pérez-González et al. (2019); Dias and Marques (2022); An et al. (2018); Hsu and Chen (2020). In this context, the interplay between long-range hopping elements and non-Hermiticity is still in its infancy as far as HOT systems are concerned  Liu et al. (2019); Ghosh and Nag (2022); Liu et al. (2021, 2023); Mondal and Nag . Therefore, considering a two-dimensional (2D) long-range NH SOTI, we, therefore, examine whether non-Hermiticity induces exceptional SOTI phases, otherwise absent in the Hermitian case, and address the following interesting questions that have not been explored so far in literature, to the best of our knowledge: How does the BBC change in the above NH phases? Can we characterize these emerging exceptional topological phases by bi-orthogonalized non-Bloch winding number?
We consider a CS-preserved generic model, hosting SOTI phases in the presence of non-Hermiticity and long-range hopping terms. The first [second] nearest neighbor (NN) Hermitian model Hamiltonian can host four [four and sixteen] zero-energy corner modes while exceptional points, caused by the NH effects, reshape topological phase boundaries as compared to the Hermitian case. As a result, we find that BBC is not only different from their Hermitian counterparts but also non-trivially modified once long-range terms are included. The phase boundaries between four [one] and one [none] zero-energy modes per corner in the second [first] NN model are revised following the dressed mass term due to the non-Bloch nature of wavefunction for NH Hamiltonian. We compute the quadrupolar winding number (QWN) appropriately in real space by exploiting the CS as well as implementing the bi-orthogonalization and non-Bloch nature to lay out the phase diagram which is in accordance with the complex spectrum under open boundary conditions (OBC).
The remainder of the article is organized in the following way. We discuss the details of the tight-binding Hamiltonian in Sec. II, where both the NH short-range and relatively long-range models are demonstrated. Sec. III is devoted to the main results of this article. In particular, we discuss the result associated with the first and the second NN neighbor hopping models in Secs. III.1 and III.2, respectively. Next, we illustrate the exceptional phase diagram in Sec. III.3 by examining the QWN. We finally summarize and conclude in Sec. IV.
II Model Hamiltonian
We consider the SOTI model in the presence of the second NN hopping as follows Ghosh and Nag (2022)
$H_{0}(\boldsymbol{\mathrm{k}})$ $=\left(\lambda^{s}_{1}\sin k_{x}+\lambda^{s}_{2}\sin 2k_{x}\right)\Gamma_{1}+\!\left(\lambda^{s}_{1}\sin k_{y}\!+\!\lambda^{s}_{2}\sin 2k_{y}\right)\Gamma_{2}\!+\!\left[m_{0}-\lambda^{h}_{1}\left(\cos k_{x}+\cos k_{y}\right)\!-\!\lambda^{h}_{2}\left(\cos 2k_{x}+\cos 2k_{y}\right)\right]\!\Gamma_{3}$
$+\left[\lambda^{f}_{1}(\cos k_{x}-\cos k_{y})+\lambda^{f}_{2}(\cos 2k_{x}-\cos 2k_{y})\right]\Gamma_{4}\ ,$ (1)
where, $\Gamma_{1}=\sigma_{x}s_{z}$, $\Gamma_{2}=\sigma_{y}s_{0}$, $\Gamma_{3}=\sigma_{z}s_{0}$, $\Gamma_{4}=\sigma_{x}s_{x}$. We consider strengths of first (second) NN hopping, spin-orbit coupling, and $C_{4}$ symmetry breaking mass terms as $\lambda^{h}_{1}$, $\lambda^{s}_{1}$, and $\lambda^{f}_{1}$, ($\lambda^{h}_{2}$, $\lambda^{s}_{2}$, and $\lambda^{f}_{2}$) respectively. Here, $m_{0}$ is the staggered mass term. In what follows, we refer to the case of $\lambda_{1}\neq 0$ and $\lambda_{2}=0$ ($\lambda_{1,2}\neq 0$) as first (second) NN model. The SOTI phase with four zero-energy corner modes arises for $-2\lambda^{h}_{1}<m_{0}<2\lambda^{h}_{1}$ when $(\lambda^{f}_{1},\lambda^{f}_{2}=0)$. The SOTI phases with four and sixteen zero-energy corner modes exist for $-2\lambda^{h}_{2}<m_{0}<0$ and $0<m_{0}<2\lambda^{h}_{1}+2\lambda^{h}_{2}$, respectively when $\lambda^{f}_{1,2}\neq 0$. Note that, this model also hosts first order topological insulator phase for $-2\lambda^{h}_{1}<m_{0}<2\lambda^{h}_{1}$ and $-2\lambda^{h}_{2}<m_{0}<2\lambda^{h}_{1}+2\lambda^{h}_{2}$, respectively, when $(\lambda^{f}_{1},\lambda^{f}_{2})=(\neq 0,0)$ and (0,0)($\neq$ 0$\neq$ 0)( ≠ 0 , ≠ 0 ). Interestingly, there is a topological phase transition that occurs at $m_{0}=0$ for the second NN model only. The above model preserves particle-hole symmetry (PHS) $P=A\mathcal{K}$ and CS $C=\sigma_{x}s_{y}$ such that $PH_{0}(\boldsymbol{\mathrm{k}})P^{-1}=-H_{0}(-\boldsymbol{\mathrm{k}})$ and $CH_{0}(\boldsymbol{\mathrm{k}})C^{-1}=-H_{0}(\boldsymbol{\mathrm{k}})$ with $A=\sigma_{x}s_{z}$ and $\mathcal{K}$ being the complex-conjugation. Importantly, the time-reversal symmetry $T=B\mathcal{K}$ is broken when $\lambda^{f}_{1,2}\neq 0$ such that $TH_{0}(\boldsymbol{\mathrm{k}})T^{-1}\neq H_{0}(-\boldsymbol{\mathrm{k}})$ with $B=\sigma_{0}s_{y}$.
Having demonstrated the physics of Hermitian SOTI model, we now focus on the NH version of the above. We associate the NH effect to the spin-orbit coupling part of the above Hamiltonian: $H_{\gamma}(\boldsymbol{\mathrm{k}})=H_{0}(\boldsymbol{\mathrm{k}})+i\gamma(\Gamma_{1}+\Gamma_{2})$ resulting in $H^{\dagger}_{\gamma}(\boldsymbol{\mathrm{k}})\neq H_{\gamma}(\boldsymbol{\mathrm{k}})$. The non-Hermiticity considered here can be thought of as an imaginary fictitious Zeeman field Yao et al. (2018); Ghosh and Nag (2022). Notice that $H_{\gamma}(\boldsymbol{\mathrm{k}})$ preserves PHS, generated by $A$ as follows $AH^{*}_{\gamma}(\boldsymbol{\mathrm{k}})A^{-1}=-H_{\gamma}(-\boldsymbol{\mathrm{k}})$. The CS continues to be preserved as $CH_{\gamma}(\boldsymbol{\mathrm{k}})C^{-1}=-H_{\gamma}(\boldsymbol{\mathrm{k}})$. We exploit the CS to define the NH analog of quadrupole moment in real space (see latter text for discussion). In the rest of the paper, we consider $\lambda^{h}_{1,2}=1$ for the sake of simplicity.
Figure 1: The energy spectra under PBC and OBC are illustrated as function of the staggered mass term $m_{0}$ in upper and lower panels, respectively, for first NN hopping. The above panels correspond to (a,d) Re[$E$], (b,e) Im[$E$] and (c,f) $|E|$. The model parameters are chosen as $\lambda^{s}_{1}=\lambda^{h}_{1}=\lambda^{f}_{1}=1.0$, and $\gamma=0.2$. Here, green lines correspond to the exceptional points obtained from the PBC case $m_{0}=\pm 2\pm\tilde{\gamma}$. The red lines, representing the exceptional phase boundary under OBC, are given by $m_{0}=\pm(2+\gamma^{2}_{1})$.
III Results
In this section, we discuss the main results of this manuscript. We analyze the eigenvalue spectra and local density of states (LDOS) corresponding to our NH model with first and second NN hopping. Afterward, we define the quadrupolar winding number for our system and demonstrate the phase diagram.
III.1 NH Model with first NN hopping
We begin with the energy dispersion of the NH SOTI model in the presence of the first NN, as shown in Figs. 1 (a,b,c) [(d,e,f)] under periodic boundary condition (PBC) [(OBC)]. While employing PBC, we find ${\rm Re}[E]=0$ for $2-\tilde{\gamma}<m_{0}<2+\tilde{\gamma}$ and $-2-\tilde{\gamma}<m_{0}<-2+\tilde{\gamma}$ with $\tilde{\gamma}=\sqrt{2\gamma^{2}}$ around $m_{0}=\mp 2$ as depicted by the gapless regions in Fig. 1(a), bounded by the green lines. These exceptional boundaries around $m_{0}=2$, $-2$, respectively, can be understood from the two-fold degeneracies of energy bands at $\boldsymbol{\mathrm{k}}=(0,0)$ and $(\pi,\pi)$ such that $\lvert E(\boldsymbol{\mathrm{k}}_{\rm EP})\rvert=0$. Interestingly, these bulk gapless exceptional points $m_{0}=\pm 2\pm\tilde{\gamma}$ are exclusively noticed in $|E|$ under the PBC case as depicted in Fig. 1(c) by the green lines.
[…]
We find that the complex energy spectra obtained under OBC and depicted in Figs. 1(d,e,f) do not mimic the underlying PBC nature. One can observe SOTI modes for which the real part of energy vanishes according to $-2-\gamma_{1}^{2}<m_{0}<2+\gamma_{1}^{2}$ with $\gamma_{1}=\gamma/\lambda^{s}_{1}$ as depicted by the red lines in Fig. 1(d). These boundaries can be anticipated by the non-Bloch form of momentum $k_{i}\to k^{\prime}_{i}-i\gamma/\lambda^{s}_{1}$ with $i=x,y$ where the renormalized mass term $m^{\prime}_{0}=m_{0}-2-\gamma_{1}^{2}<0~{}(>0)$ determines the topological (trivial) phase of the NH model Yao et al. (2018); Ghosh and Nag (2022). Note that, for Bloch momentum $\boldsymbol{\mathrm{k}}=(0,0)$ [$\boldsymbol{\mathrm{k}}=(\pi,\pi)$], the exceptional phase boundaries extend till $m_{0}=\pm(2+\gamma_{1}^{2})$ leading to the emergence of exceptional SOTI phases beyond the Hermitian gapless phase boundaries $m_{0}=\pm 2$. All the single-particle energy states under OBC except the corner modes exhibit an imaginary component of energy for $|m_{0}|<2$ as shown in Fig. 1(e). This is markedly different from the PBC case, depicted in Fig. 1(b), where single particle states have finite amount of imaginary energy for $|m_{0}|>2$. This refers to a macroscopic degeneracy within a certain range of $m_{0}$ as far as the ${\rm Im}[E]$ is considered. Since such macroscopic degeneracy does not exist for ${\rm Re}[E]$, the $|E|$ demonstrates the NH corner modes for $|m_{0}|<|2+\gamma_{1}^{2}|$ under OBC [see red lines in Fig. 1(f)], while non-Hermiticity mediated bulk gapless points are noticed for the PBC case, see Fig. 1(c).
[…]
III.2 NH Model with second NN hopping
Figure 3: The energy spectra under PBC and OBC are illustrated in upper and lower panels, respectively as a function of $m_{0}$, for second NN hopping. The panels correspond to (a,d) Re[$E$] (b,e) Im[$E$] and (c,f) $|E|$. We choose the model parameters as $\lambda^{s}_{1}=\lambda^{h}_{1}=\lambda^{f}_{1}=1.0$, $\lambda^{s}_{2}=\lambda^{h}_{2}=\lambda^{f}_{2}=1.0$ and $\gamma=0.2$. The green lines correspond to the exceptional points $m_{0}=4\pm\tilde{\gamma}$, ±$\pm\tilde{\gamma}$, and $-2\pm\tilde{\gamma}$ under PBC. The red lines, representing the exceptional phase boundary under OBC, are given by $m_{0}=4+5\gamma_{2}^{2}$, $3\gamma_{2}^{2}$, and $-2-\gamma_{2}^{2}$.
To start with, we depict the energy dispersion of the NH SOTI model in the presence of the second NN in Figs. 3 (a,b,c) [(d,e,f)] under PBC [OBC]. We find degenerate eigenstate with ${\rm Re}[E]=0$ under PBC for $-2-\tilde{\gamma}<m_{0}<-2+\tilde{\gamma}$, $-\tilde{\gamma}<m_{0}<\tilde{\gamma}$, and $4-\tilde{\gamma}<m_{0}<4+\tilde{\gamma}$ as depicted by the green lines in Fig. 3(a). These exceptional boundaries around $m_{0}=-2$, 00 and $4$, respectively, can be understood from the two-fold degeneracies of energy bands at $\boldsymbol{\mathrm{k}}=(\pm 2\pi/3,\pm 2\pi/3)$, $(\pi,\pi)$, and $(0,0)$ such that $\lvert E(\boldsymbol{\mathrm{k}}_{\rm EP})\rvert=0$. Similar to the earlier first NN model, these bulk gapless exceptional points $m_{0}=-2\pm\tilde{\gamma}$, ±$\pm\tilde{\gamma}$, and $4\pm\tilde{\gamma}$ are exclusively observed in $|E|$ for PBC case as depicted in Fig. 3(c) by green lines.
We now examine the energy spectrum for the second NN model in Figs. 3(d,e,f) employing OBC. Similar to the first NN case, the momentum takes the following non-Bloch form $k_{i}\to k^{\prime}_{i}-i\gamma_{2}$ with $\gamma_{2}=\gamma/(\lambda^{s}_{1}+2\lambda^{s}_{2})$ and $i=x,y$. This leads to the renormalized mass term $m^{\prime}_{0}=m_{0}-4-5\gamma_{2}^{2}<0~{}(>0)$ for the topological (trivial) phase with zero-energy (finite energy bulk) modes considering the Bloch momentum $\boldsymbol{\mathrm{k}}=(0,0)$. On the other hand, another topological (trivial) phase with zero-energy (finite energy) modes appears for $-m^{\prime}_{0}=m_{0}+2+\gamma_{2}^{2}>0~{}(<0)$ while exploiting energy around the Bloch momentum $\boldsymbol{\mathrm{k}}=(\pm 2\pi/3,\pm 2\pi/3)$. Our analysis indicates the existence of four [sixteen] corner modes for $3\gamma_{2}^{2}<m_{0}<4+5\gamma_{2}^{2}$ [$-2-\gamma^{2}_{2}<m_{0}<3\gamma_{2}^{2}$] yielding ${\rm Re}[E]=0$ under OBC. However, the numerical findings shown in Fig. 3(d) and the topological regime highlighted by the red lines, do not fully match with the exceptional boundaries predicted analytically. This can be attributed to the finite size effect in the second NN case, which is substantially small for the first NN case. We find macroscopic degeneracies at ${\rm Re}[E]=0$ around $m_{0}=0$, unlike the previous case. The complex energy spectrum ${\rm Im}[E]$ does not manifest any noteworthy features, as shown in Figs. 3(b,e), irrespective of PBC and OBC cases. The absolute value of energy is expected to vanish i.e.,  $|E|=0$ under OBC for $-2-\gamma^{2}_{2}<m_{0}<4+5\gamma_{2}^{2}$. However, numerical results suffer from finite size effects as far as the exceptional boundaries are concerned [see the red lines in Fig. 3(f)]. The finite value of $\gamma$ (NH effect) thus extends the Hermitian topological phase beyond its boundaries, leading to exceptional topological phases.
[…]
III.3 Quadrupole Winding number
We now investigate the topological invariant, namely, quadrupole winding number (QWN), by exploiting the CS. In the following discussion, we first illustrate the Hermitian version of QWN. Given the fact that CS constraints $CH_{0}(\boldsymbol{\mathrm{k}})C^{-1}=-H_{0}(\boldsymbol{\mathrm{k}})$, we can anti-diagonalize the Hamiltonian in the basis of the CS operator spanned by $U_{C}$ as follows  Ryu et al. (2010); Chiu et al. (2016)
$\tilde{H}_{0}=U_{C}^{\dagger}H_{0}U_{C}=\begin{pmatrix}0&h\\ \tilde{h}&0\end{pmatrix}\ ,$ (2)
Here, $\tilde{h}=h^{\dagger}$ if $\tilde{H}_{0}$ is Hermitian. We find $U_{C}CU^{\dagger}_{C}=\pm 1$, suggesting that CS can be classified into two kinds of sub-lattices, namely $A$ and $B$ for $+$ and -- expectation values, respectively. This further entails that $U_{C}=U^{A}_{C}-U^{B}_{C}$ where $U^{A}_{C}=\sum_{\alpha\in A}|\alpha\rangle\langle\alpha|$ and $U^{B}_{C}=\sum_{\beta\in B}|\beta\rangle\langle\beta|$.
Employing singular value decomposition of $h$, we obtain $h=U_{A}\Sigma U_{B}^{\dagger}$ where $U_{A,B}$ are unitary matrices and $\Sigma$ denotes a diagonal matrix. Note that, the diagonal elements of $\Sigma$ are referred to as singular values. One can compute the flattened Hamiltonian $Q$, having eigenvalue ±$\pm 1$ as follows Benalcazar and Cerjan (2022)
$Q=\begin{pmatrix}0&q\\ q^{\dagger}&0\end{pmatrix}\ ,$ (3)
with $q=U_{A}U^{\dagger}_{B}$ being a unitary matrix. It has been shown that the winding number, derived using $q$ and $q^{\dagger}$, is related to the relative polarization of $A$ and $B$ sublattice. In a similar spirit, the winding number in the real space is given by Lin et al. (2021)
$\nu=\frac{1}{{2\pi i}}{\rm{Tr}}\left[{\log\left({{{\mathcal{X}}_{A}}\mathcal{X}_{B}^{\dagger}}\right)}\right]\ ,$ (4)
where, ${{\mathcal{X}}_{\sigma}}=U_{\sigma}^{\dagger}U^{\sigma}_{C}\mathcal{X}U^{\sigma}_{C}{U_{\sigma}}$ ($\sigma=A,B$) are unitary matrices. The operator ${{\mathcal{X}}_{\sigma}}$ denotes the sublattice dipole operator, which is the projection of the position operator onto the $\sigma$ sector of the chiral basis. The position operator, i.e.,  the dipole operator $\mathcal{X}=\exp(2i\pi x/L)$ is defined on a periodic array of one-dimensional length $L$.
Now turning to the two-dimensional system where the dipole operator $\mathcal{X}$ can be replaced by the quadrupole operator $\mathcal{Q}=\exp(2i\pi xy/L_{x}L_{y})$. This results in the sublattice quadrupole operator $\mathcal{Q}_{\sigma}=U_{\sigma}^{\dagger}U^{\sigma}_{C}\mathcal{Q}U^{\sigma}_{C}U_{\sigma}$. Therefore, the QWN can be defined as Benalcazar and Cerjan (2022)
$N_{xy}=\frac{1}{2\pi i}{\rm Tr}\left[\log\left(\mathcal{Q}_{A}\mathcal{Q}^{\dagger}_{B}\right)\right]\ ,$ (5)
This invariant is quantized to an integer number and predicts the number of topologically protected corner states at each corner of the 2D lattice.
We now examine the present situation with $\gamma\neq 0$ where the NH analog of QWN is discussed. Importantly, CS is also preserved for the NH Hamiltonian $CH_{\gamma}(\boldsymbol{\mathrm{k}})C^{-1}=-H_{\gamma}(\boldsymbol{\mathrm{k}})$ allowing for the anti-diagonal form of $\tilde{H}_{\gamma}$. At the same time, the definition of $U^{A,B}_{C}$ remains unaltered for the NH case. We adopt the bi-orthogonalized definitions of $U_{A}$ and $U^{\dagger}_{B}$, obtained from the singular value decomposition of $h$, to define $U^{\dagger}_{A}$ and $U_{B}$, respectively. One has to ensure $\sum_{n}|U^{R}_{\sigma,n}\rangle\langle U^{L}_{\sigma,n}|=\mathds{1}$ and $\langle U^{L}_{\sigma,n}|U^{R}_{\sigma,m}\rangle=\delta_{mn}$ with $\sigma=A,B$ and $L(R)$ denotes the left (right) eigenvectors. This results in left [right] eigenvector corresponding to right eigenvector $(U^{\dagger}_{B})^{\dagger}\equiv U^{R}_{B}$ [left eigenvector $U^{\dagger}_{A}\equiv U^{L}_{A}$] as $U^{\dagger}_{B}\equiv U^{L}_{B}$ [$U_{A}\equiv U^{R}_{A}$]. Therefore, sublattice quadrupole operator $\mathcal{Q}_{\sigma}$ takes the form $\mathcal{Q}_{\sigma}=U_{\sigma}^{L}U^{\sigma}_{C}\mathcal{Q}U^{\sigma}_{C}U^{R}_{\sigma}$. In addition, the non-Bloch form of momentum has to be incorporated while computing $\mathcal{Q}_{\sigma}$.
To be precise, the complex momentum $k_{i}\to k^{\prime}_{i}-i\gamma_{2}$ with $i=x,y$ leads to the exponentially enhanced and suppressed hopping elements by the multiplicative factors $\exp(\gamma_{1})$ and $\exp(-\gamma_{1})$ [$\exp(\gamma_{2})$ and $\exp(-\gamma_{2})$] for first [second] NN models. We use the real space form of the tight-binding model with the renormalized hopping amplitudes as follows $\lambda^{s,h,f}_{1,2}\to\lambda^{s,h,f}_{1,2}\exp(\gamma_{1,2})$ and $\lambda^{s,h,f}_{1,2}\to\lambda^{s,h,f}_{1,2}\exp(-\gamma_{1,2})$ for forward and backward hopping amplitudes, respectively. We consider the real space version of NH Hamiltonian $H_{\gamma}$ with the above mentioned renormalized hoppings in order to compute QWN. Altogether this enables us to define the NH analog of QWN $N_{xy}$ with dressed hopping and bi-orthogonalized definition. Note that, the real part of $N_{xy}$ exhibits quantized value for the present case with $\gamma\neq 0$ as demonstrated below.
We now discuss the phase diagram of the first and second NN NH model in the $\gamma$-$m_{0}$ plane in Figs. 5(a,b) respectively. Note that, there exists four corner modes with ${\rm Re}[E]=0$ for $m_{0}<2+\gamma_{1}^{2}$, yielding $N_{xy}=1$. This refers to the fact that there exists only one topological zero-mode per corner [see Fig. 5(a)]. On the other hand, when $m_{0}>2+\gamma_{1}^{2}$, the NH model does not host any topological phase and hence $N_{xy}=0$. Hence, the topological phase boundary is $m_{0}=2+\gamma_{1}^{2}$ which is indicated by the yellow solid line. This is also predicted from the complex energy spectrum with OBC. The yellow dashed lines in Fig. 5(a) represent $m_{0}=2\pm\tilde{\gamma}$ lines as predicted from the complex energy under PBC. Interestingly, the real space invariant QWN fails to identify these phase boundaries. Unlike the Hermitian system, the phase boundaries between topological and trivial phases cannot be captured by the energy spectrum under OBC and PBC in the present NH case. This clearly suggests that the topological phase, predicted from the complex energy spectrum in OBC, is apprehended by the non-Bloch and bi-orthogonalized version of QWN. We note that the analytically derived phase boundaries are valid for $\gamma_{1,2}\ll 1$.
We find qualitatively similar results in case of the second NN model as shown in Fig. 5(b). In addition to $N_{xy}=1$ phase, we obtain $N_{xy}=4$ phase where four zero-energy modes with ${\rm Re}[E]=0$ are present at each corner. While investigating the phase boundaries, it is expected to find $N_{xy}=1$ [$4$] for $3\gamma_{2}^{2}<m_{0}<4+5\gamma_{2}^{2}$ [$-2-\gamma^{2}_{2}<m_{0}<3\gamma_{2}^{2}$]. For smaller strength of non-Hermiticity i.e.,  $\gamma\to 0$, we find quantitative agreement between the analytical and numerical findings. The phase boundaries $m_{0}=4+5\gamma_{2}^{2}$, $3\gamma_{2}^{2}$, and $-2-\gamma^{2}_{2}$, designated by the yellow solid line, do not fully comply with the $N_{xy}$ profile for $\gamma>0.1$. This can be due to more intricacies than just the finite size effect. Interestingly, the following tendency is noticed: for $\tilde{\gamma}<m_{0}<4+5\gamma_{2}^{2}$, one obtains $N_{xy}=1$ while within the regime $-2-\tilde{\gamma}<m_{0}<-\tilde{\gamma}$, $N_{xy}$ acquires the value $4$. Therefore, the non-Bloch and bi-orthogonalized version of QWN can quantitatively and qualitatively identify the phase boundaries of SOTI phases hosting four and sixteen corner modes starting from the trivial phases for $m_{0}>4+5\gamma_{2}^{2}$ and $m_{0}<-2-\tilde{\gamma}$, respectively, across which $N_{xy}$ jumps between zero and finite values. The real space invariant QWN is thus a useful topological marker to identify the exceptional phases for our NH system with OBC.
On the other hand, between two SOTI phases with different number of corner modes (for $-\tilde{\gamma}<m_{0}<\tilde{\gamma}$), we find $N_{xy}\neq 0$ as depicted in Fig. 5(b). In the complex energy analysis under OBC, we find macroscopic degeneracies with ${\rm Re}[E]=0$ for $-2-\gamma^{2}_{2}<m_{0}<\tilde{\gamma}$. Likewise the earlier first NN case, $N_{xy}$ does not exhibit any jumps between finite and zero values around $m_{0}=4\pm\tilde{\gamma}$ and $m_{0}=-2\pm\tilde{\gamma}$ which are predicted by the complex energy spectrum under PBC. On the contrary, $m_{0}=\pm\tilde{\gamma}$ boundaries, predicted by complex energy spectrum under PBC, are visible as $N_{xy}$ changes between two finite values. The exceptional lines $m_{0}=-2\pm\tilde{\gamma}$, ±$\pm\tilde{\gamma}$, and $4\pm\tilde{\gamma}$ obtained employing PBC are depicted by yellow dashed lines in Fig. 5(b). This agreement is surprising and yet to be explored in the future. However, there is an apparent discrepancy between the solid yellow lines and the numerically obtained $N_{xy}$. This mismatch is owing to the fact that the mathematical form of non-Bloch transformation that we consider in the hopping terms while computing $N_{xy}$, employing PBC, is computed by employing a low-energy version of $H_{\gamma}(\boldsymbol{\mathrm{k}})$. To obtain the low-energy spectrum of $H_{\gamma}(\boldsymbol{\mathrm{k}})$, we expand the Hamiltonian around $\boldsymbol{\mathrm{k}}=(0,0)$. By doing that we can obtain the phase boundary associated with the right part of Fig. 5(b). However, when we incorporate the higher-order hopping elements, the low-energy model around $\boldsymbol{\mathrm{k}}=(0,0)$ does not necessarily encapsulate all the phase transition lines. In that scenarion, one should also consider a low-energy Hamiltonian around other momenta such as $\boldsymbol{\mathrm{k}}=(\pm 2\pi/3,\pm 2\pi/3)$, depending upon the value of $m_{0}$. Nevertheless, this scenario adds substantial complexity to the problem as one should consider a different non-Bloch form for different $m_{0}$. Thus, finding a universal transformation to obtain the exact phase boundary in the NH second NN hopping case still remains an interesting question and beyond the scope of the present paper. Nevertheless, complex energy spectra under PBC might be useful for understanding the phase boundaries between two different topological phases.
Figure 5: The phase diagram in $m_{0}$-$\gamma$ plane for NH first NN and second NN hopping models are shown in panel (a), and (b), respectively. The color bar represents QWN $N_{xy}$. The phase boundary $m_{0}=2+\gamma_{1}^{2}$ [$m_{0}=2\pm\tilde{\gamma}$], obtained from OBC [PBC], between exceptional topological and trivial phases in panel (a) is identified by the yellow solid [dashed] line. The phase boundaries, associated with OBC [PBC], between topological phases hosting sixteen and four corner modes, and trivial phase are given by $m_{0}=4+5\gamma_{2}^{2}$ [$m_{0}=4\pm\tilde{\gamma}$], $m_{0}=3\gamma_{2}^{2}$ [$m_{0}=\pm\tilde{\gamma}$], and $m_{0}=-2-\gamma_{2}^{2}$ [$m_{0}=-2\pm\tilde{\gamma}$], respectively, represented by the yellow solid [dashed] lines in panel (b).
IV Summary and Conclusion
To summarize, in this article, we consider a long-range hopping model in the presence of non-Hermiticity to investigate the emergence of second-order topological phases. By exploring the real part of the complex energy spectrum for the first and second NN NH models under OBC, we find that the former model only hosts four zero-energy corner modes while the latter model can host four as well as sixteen zero-energy corner modes as the hallmark of the NH SOTI phases. We compute the real space invariant namely, bi-orthogonalized QWN, by keeping in mind the non-Bloch form of the momentum to uniquely characterize the different topological phases. The phase boundaries, captured by the above invariant, can successfully mimic the emergence of NH SOTI phases out of the trivial phases, as demonstrated by the complex energy dispersion under OBC for both the first and second NN models. The topological phase boundary between two different topological phases, observed in the second NN model, can be anticipated from the complex energy spectrum under PBC for the above model. In the future, one can include disorder to study the exceptional topological Anderson insulators hosting corner modes where a generalized version of the presently adopted real space topological index will have to be examined.
Acknowledgments
A.K.G. and A.S. acknowledge SAMKHYA: High-Performance Computing Facility provided by Institute of Physics, Bhubaneswar, for numerical computations. TN acknowledges the NFSG “NFSG/HYD/2023/H0911” from BITS Pilani.